# From Unaligned to Aligned: Scaling Multilingual LLMs with Multi-Way Parallel Corpora

### v1

@@ Abstract
Yingli Shen1thanks: Equal contribution., Wen Lai2,3∗, Shuo Wang1 Kangyang Luo1, Alexander Fraser2,3, Maosong Sun1thanks: Corresponding author. 1 Tsinghua University, China 2 Technical University of Munich, Germany 3 Munich Center for Machine Learning, Germany syl@mail.tsinghua.edu.cn, wen.lai@tum.de
@@ 3 Experimental Setup
To isolate the effects of multi-way parallel data, we conduct both continued pretraining (Parmar et al., 2024) and instruction tuning (Zhang et al., 2023) on TED2025  [^4] For RQ1 and RQ2, we focus on continued pretraining to avoid interference from instruction tuning. For RQ3, we instead focus on instruction tuning to assess its specific impact.. We experiment with two model families and sizes: LLaMA-3.1-8B / LLaMA-3.1-8B-Instruct and Qwen-2.5-14B / Qwen-2.5-14B-Instruct (available on Hugging Face [^5] https://huggingface.co). To make fine-tuning feasible, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2022) instead of performing full parameter updates. Full hyperparameter settings and training configurations are provided in Appendix B.
@@ 4 Effectiveness of Multi-Way Corpora
For fairness, we fix the total continued pretraining data at 5 million tokens (5M) and evaluate two LLM backbones (LLaMA-3-8B and Qwen-2.5-14B) under three conditions: (1) Multi-Way: Pretraining on our multi-way parallel corpus (TED2025); (2) Unaligned: Pretraining on an unaligned multilingual corpus (DCAD-2000; Shen et al., 2025); (3) Baseline: The original pretrained checkpoint without additional data.
@@ 4.1 Downstream Performance
We evaluate all variants in a zero-shot setting across four benchmarks—MMMLU, XCOPA, FLORES-101, and xIFEval—that cover understanding, reasoning, generation, and instruction following. The full results are reported in Table 3. Our findings show that, across all tasks, Multi-Way consistently outperforms both Baseline and Unaligned across both low- and high-resource languages. For example, on MMMLU, Multi-Way achieves accuracies of 22.48/41.38 in low/high-resource languages, compared to 18.27/33.72 for the Baseline and 19.64/36.26 for Unaligned. Similar improvements are observed on FLORES-101 and xIFEval. These results demonstrate that multi-way alignment provides stronger cross-lingual supervision, thereby enhancing both discriminative and generative capabilities.
@@ 4.3 Cross-Lingual Representation Alignment
Table 4 shows that Multi-Way outperforms the other models, yielding higher CKA and better retrieval accuracy. Figure 4 further corroborates these results: Multi-Way demonstrates denser SVCCA alignments, particularly for linguistically distant language pairs. These metrics confirm that multi-way pretraining promotes a more coherent, language-agnostic embedding space, which drives the observed improvements in downstream performance and cross-lingual transfer.
@@ 5 Impact Factors
In Section 4, we demonstrate that using multi-way parallel data can significantly enhance the multilingual capabilities of LLMs. To better understand the factors driving this improvement, we analyze two key aspects [^6] Additional analyses on data size and language-family combinations can be found in Appendix C., while keeping the total pretraining fixed at 5 million tokens: (1) Degree of parallelism: the number of languages aligned in each training example (Section 5.1). (2) English as a pivot: the impact of including versus excluding English in multi-way groups (Section 5.2).
@@ 5.2 English as Pivot
English is widely used as a pivot language in multilingual MT and NLP Kim et al. (2019); Mallinson et al. (2018); Lai et al. (2023b). To explore its impact, we form five groups [^7] We provide the configurations of each group in Table 8 of Appendix C.2, each with both English-included and English-excluded variants across six languages. Figure 6 compares their performance across various tasks.
@@
Figure 6: Performance (%) comparison of models with and without (w/o) English across five different language groupings.
@@ 6 Instruction Tuning
In Sections 4 and 5, we demonstrate that using multi-way data can significantly improve multilingual performance. In this section, we further investigate whether instruction tuning can also enhance multilingual performance effectively. Specifically, we address the following key questions [^8] Additionally, we explore the cross-lingual alignment of representations in instruction tuned models. Detailed experimental settings and results can be found in the Appendix D.: (1) Which of the different instruction fine-tuning objectives, built on multi-way parallel data, is most effective? (2) Do models trained with multi-way parallel data exhibit better generalization across domains?
@@ Appendix A Statistics of the Constructed TED2025
In Figure 2, we compare the translation quality of TED2025 with existing multi-way parallel datasets, demonstrating the overall effectiveness of TED2025 translations. To offer a more comprehensive and intuitive view of translation quality, Table 1011121314, and 15 report COMET-QE score for all 4,765 language pairs included in the dataset. This analysis highlights that, despite being a multi-way parallel corpus, TED2025 can be readily decomposed into bilingual sentence pairs, making it suitable for training machine translation models or fine-tuning LLMs on specific language pairs. By providing both the number of bilingual pairs and their associated translation quality, we aim to support researchers in selecting suitable data for their specific translation tasks and in optimizing performance on targeted language pairs.
@@
Appendix C Impact Factors: Additional Analysis
In Section 5, we investigated two key factors affecting the effectiveness of leveraging multi-way parallel data to enhance the multilingual capabilities of LLMs: the degree of parallelism and the presence of English in the language combinations. In this section, we extend our analysis to two additional factors: the size of the training data and the linguistic characteristics of the languages involved in the combinations.
@@ Appendix C Impact Factors: Additional Analysis
C.1 Training Data Size
C.2 Language Combinations
In Section 5.2, we investigated the influence of including English in sampling combinations on model performance. In this section, we extend that analysis by investigating an alternative sampling strategy: whether the selected languages belong to the same language family. To this end, we construct four language groups. Groups 1 and 2 consist of languages from the same language family, whereas Groups 3 and 4 include languages from diverse families. The specific language configurations for each group are detailed in Table 8.
@@ C.2 Language Combinations
Appendix D Representation Alignment with Instruction Tuning
@@
art 5858 gender 144 product design 68 Audacious Project 35 television 17

### v4

@@ Abstract
Yingli Shen1thanks: Equal contribution., Wen Lai2,3∗, Shuo Wang1, Ge Gao4 Kangyang Luo1, Alexander Fraser2,3, Maosong Sun1,5thanks: Corresponding author. 1 Department of Computer Science and Technology, Tsinghua University 2 Technical University of Munich  3 Munich Center for Machine Learning 4 Minzu University of China  5 Institute for AI, Tsinghua University syl@mail.tsinghua.edu.cn, wen.lai@tum.de
@@ 3 Experimental Setup
To isolate the effects of multi-way parallel data, we conduct both continued pretraining (Parmar et al., 2024) and instruction tuning (Zhang et al., 2023) on TED2025. We experiment with two model families and sizes: LLaMA-3.1-8B / LLaMA-3.1-8B-Instruct and Qwen-2.5-14B / Qwen-2.5-14B-Instruct (available on Hugging Face [^4] https://huggingface.co). To make fine-tuning feasible, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2022) instead of performing full parameter updates. Full hyperparameter settings and training configurations are provided in Appendix B.
@@ 4 Effectiveness of Multi-Way Corpora
For fairness, we fix the total continued pretraining data at 5 million tokens (5M) and evaluate two LLM backbones (LLaMA-3-8B and Qwen-2.5-14B) under three conditions: (1) Multi-Way: Pretraining on our multi-way parallel corpus (TED2025); (2) Unaligned: Pretraining on an unaligned multilingual corpus [^5] Unaligned data is sourced from DCAD-2000 Shen et al. (2025) rather than TED2025, due to considerations related to scalability and domain coverage. (DCAD-2000; Shen et al., 2025); (3) Baseline: The original pretrained checkpoint without additional data.
@@ 4.1 Downstream Performance
We evaluate all variants in a zero-shot setting across four benchmarks—MMMLU, XCOPA, FLORES-101, and xIFEval—that cover understanding, reasoning, generation, and instruction following. The full results are reported in Table 3. Our findings show that, across all tasks, Multi-Way consistently outperforms both Baseline and Unaligned across both low- and high-resource languages. For example, on MMMLU, Multi-Way achieves accuracies of 22.48/41.38 in low/high-resource languages, compared to 18.27/33.72 for the Baseline and 19.64/36.26 for Unaligned. Similar improvements are observed on FLORES-101 and xIFEval. These results demonstrate that multi-way alignment provides stronger cross-lingual supervision, thereby enhancing both discriminative and generative capabilities. Although our main experiments focus on English–X translation directions, we also evaluate the trained models on non-English-centric translation pairs. The results, which show that models trained on aligned data significantly outperform those trained on unaligned data even for non-English-centric tasks, are provided in Appendix C.
@@ 4.3 Cross-Lingual Representation Alignment
Table 4 shows that Multi‑Way outperforms the other models, yielding higher CKA and better retrieval accuracy. Figure 4 further corroborates these results: Multi‑Way demonstrates denser SVCCA alignments, particularly for linguistically distant language pairs. These metrics confirm that multi-way pretraining promotes a more coherent, language-agnostic embedding space, which drives the observed improvements in downstream performance and cross-lingual transfer.
@@ 5 Impact Factors
In Section 4, we demonstrate that using multi-way parallel data can significantly enhance the multilingual capabilities of LLMs. To better understand the factors driving this improvement, we analyze two key aspects, while keeping the total pretraining fixed at 5 million tokens: (1) Degree of parallelism: the number of languages aligned in each training example (Section 5.1). (2) English as a pivot: the impact of including versus excluding English in multi-way groups (Section 5.2). (3) Language combinations: the role of language family composition in training data (Section 5.3). (4) Training data size: how the amount of pretraining data affects model performance (Section 5.4).
@@ 5.2 English as Pivot
English is widely used as a pivot language in multilingual MT and NLP Kim et al. (2019); Mallinson et al. (2018); Lai et al. (2023b). To explore its impact, we form five groups [^6] We provide the configurations of each group in Table 10 of Appendix D, each with both English-included and English-excluded variants across ten languages. Table 5 compares their performance across various tasks.
@@
MMMLU XCOPA
FLORES
(Eng-X)
FLORES
(X-Eng)
xIFEval
Group 1 with 23.17 30.93 5.80 13.75 23.39
w/o 20.84 30.07 6.63 14.56 24.14
Group 2 with 22.19 35.33 6.74 13.88 22.19
w/o 18.51 31.60 7.13 14.18 22.19
Group 3 with 26.47 36.73 6.40 13.48 22.69
w/o 23.83 35.78 6.38 14.00 22.91
Group 4 with 23.42 33.84 6.20 12.11 23.66
w/o 20.26 30.84 6.67 14.76 23.67
Group 5 with 23.89 39.64 6.96 13.07 23.19
w/o 20.39 34.15 7.83 14.79 23.85
Table 5: Performance (%) comparison of models with and without (w/o) English across five different language groupings.
@@
5.3 Language Combinations
In Section 5.2, we investigate the influence of including English in sampling combinations on model performance. In this section, we extend that analysis by investigating an alternative sampling strategy: whether the selected languages belong to the same language family. To this end, we construct four language groups. Groups 1 and 2 consist of languages from the same language family, whereas Groups 3 and 4 include languages from diverse families. The specific language configurations for each group are detailed in Appendix D (Table 10).
5.4 Training Data Size
@@ 6 Instruction Tuning
In this section, we further investigate whether instruction tuning can also enhance multilingual performance effectively. Specifically, we address the following key questions [^7] Additionally, we explore the cross-lingual alignment of representations in instruction tuned models (Appendix E.): (1) Which of the different instruction fine-tuning objectives, built on multi-way parallel data, is most effective? (2) Do models trained with multi-way parallel data exhibit better generalization across domains?
@@ 7 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grants No. 62236011 and No. T2341003), and by a grant from the Guoqiang Institute, Tsinghua University. Additionally, support was received from the European Research Council (ERC) under the European Union’s Horizon Europe Research and Innovation Programme (Grant Agreement No. 101113091), as well as from the German Research Foundation (DFG; Grant FR 2829/7-1).
@@ Appendix A Statistics of the Constructed TED2025
In Figure 2, we compare the translation quality of TED2025 with existing multi-way parallel datasets, demonstrating the overall effectiveness of TED2025 translations. To offer a more comprehensive and intuitive view of translation quality, Table 1213141516, and 17 report COMET-QE score for all 4,765 language pairs included in the dataset. This analysis highlights that, despite being a multi-way parallel corpus, TED2025 can be readily decomposed into bilingual sentence pairs, making it suitable for training machine translation models or fine-tuning LLMs on specific language pairs. By providing both the number of bilingual pairs and their associated translation quality, we aim to support researchers in selecting suitable data for their specific translation tasks and in optimizing performance on targeted language pairs.
@@
sk-ta $k$-ko eo-sw ja-zh nl-tr
BLEU COMET BLEU COMET BLEU COMET BLEU COMET BLEU COMET
Baseline 3.14 51.24 3.11 50.53 2.93 48.56 5.02 54.71 3.63 52.35
Unaligned 4.16 52.75 3.82 51.97 3.46 49.73 6.39 55.84 4.75 52.27
Multi-Way 6.31 53.25 5.38 52.66 5.14 51.35 8.06 56.27 6.82 53.50
Table 9: Translation performance of models trained on aligned, unaligned, and multi-way data setups for five non-English-centric language pairs.
Appendix C Non-English-Centric Translation Performance
In addition to the primary experiments on English–X translation directions presented in Section 4.1, we also investigate the performance of our models on non-English-centric translation pairs. This analysis is motivated by the potential of multilingual models to perform well in settings where neither language involved is English. To this end, we randomly select five language pairs (sk-ta, kk-ko, eo-sw, ja-zh, and nl-tr) that do not involve English as either the source or target language. We then evaluate the translation performance of the pretrained models under three different data configurations: Baseline, Unaligned, and Multi-way, as described in Section 4.1. The results, including BLEU and COMET scores, are presented in Table 9.
We observe that: (1) Models trained on aligned data consistently outperform those trained on unaligned data across all translation pairs, indicating that aligned corpora are essential for achieving high translation quality, regardless of whether English is involved. (2) Models trained on unaligned data show a significant improvement over the baseline, suggesting that even in the absence of high-quality aligned corpora, leveraging unaligned multilingual data can still enhance translation performance. (3) The multi-way approach consistently yields superior results compared to both the baseline and unaligned models. This highlights the advantages of multi-way training for multilingual translation tasks, as it not only improves performance on English-centric tasks but also significantly boosts translation quality for non-English-centric language pairs.
@@ Appendix C Non-English-Centric Translation Performance
Appendix D Configuration of Different Language Combinations
As discussed in Section 5.3, we define four distinct language groups for our analysis. The specific configurations of these groups are presented in Table 10.
Appendix E Representation Alignment with Instruction Tuning
@@
art $58$ gender 144 product design 68 Audacious Project 35 television 17